# nsarafianos.github.io/garment3dgen

### v1

@@ 3.3 Texture Estimation
Given the untextured deformed 3D geometry $M_{\text{def}}$, we aim to generate high-fidelity textures that match the input image. We leverage a 2D text-to-image generation model to create high-quality, high-resolution textures with vivid details given a text prompt. For garment generation starting from images, a text prompt can automatically be obtained using image caption models or by specifying them manually. Our pipeline consists of the following steps: generating images of the given fashion asset from multiple views and backprojecting these images onto the mesh surface to create UV texture $T\in\mathbb{R}^{H\times W\times C}$. There are two challenges in adapting a 2D generation model to 3D objects: establishing geometry-texture correspondence and ensuring multi-view consistency.
@@
Multi-view Consistency: Directly conducting view-by-view image synthesis cannot guarantee that the generated views are consistent with each other. To solve this problem, we first synthesize the front and back views simultaneously to implicitly enforce global consistency. In addition, we leverage the fact that the geometry of fashion assets is mostly flat, and conduct depth-aware inpainting for the remaining views to ensure the newly generated textures are locally consistent. To handle occluded areas, we design an automatic view selection algorithm to inpaint the textures in a coarse-to-fine manner: from the remaining unpainted area, we select the view with the most unfilled pixels to generate the textures iteratively from large regions to small pieces.
@@
3.4 Fitting the Generated Garments to Parametric Bodies
While the aforementioned supervisions and regularizations aim to ensure that the quality of the generated garment will be satisfactory enough to be simulated, there are no guarantees that the garment shape, scale, pose and orientation will be those required to drape it on a parametric body [41] and simulate it. To accomplish this task, we run an optimization procedure during which the generated garment remains fixed in the generated pose and the pose and shape of the parametric body are transformed such that the garment can accurately fit the body. This optimization process shown in Fig. 3, starts with a rigid transformation and scaling of the body and continues with an optimization of the body pose and shape using the Chamfer distance loss. Finally, once the optimization has converged, we run an additional optimization step penalizing body-cloth collisions, ensuring that the fit is as accurate and realistic as possible.
Figure 3: 3D Garment Generation (left): Given an image or a text prompt as guidance and a base geometry mesh (bottom left inset) that can be far from the target we generate high-quality textured 3D geometries of both real as well as fantastical garments. Fitting (right): We start with the generated textured 3D garment (in this case a medieval armor) and a parametric body in its canonical pose (left) and run the body-garment optimization process to optimize for the body pose and shape parameters such that the generated garment can fit in the body accurately without penetrations.
@@ 4 Experiments
Data: We rely on a handful of artist-created garment templates covering basic cloth categories (e.g. T-shirt, shirt, tank-top, dress, etc). These untextured 3D meshes are in a canonical pose and will be publicly released in order to facilitate future 3D garment research. When it comes to images that serve as a guidance, we collect a variety of real images with garments in different poses, different textures, garments that do not exist in our mesh library and even AI generated garments both real and fantastical generated from a textual prompt.
@@
Metrics: Quantitatively evaluating our results is a challenging task in the absence of 2D or 3D ground-truth 2D for what are we trying to accomplish. However we can evaluate how consistent the geometries are to the input image and hence we render the untextured outputs of all methods from 36 views and compute their perceptual scores using the LPIPS metric [82] as well as their image-based CLIP similarity score using the cosine distance between their embeddings.
@@
Baselines: We evaluate our approach against: i) TextDeformer [18] which deforms input meshes based on text prompts, ii) ImageDeformer - a variant of TextDeformer we set up - where the input text is replaced with an image, iii) Wonder3D [40] and iv) Zero123++[60] both of which generate 3D geometries given a single image using 2D diffusion models, v) the newly introduced ZeroShape [25] which performs zero-shot reconstruction and vi) a recent triplane-based 3D Gaussian Splatting approach [84]. The first two works also take as input a base mesh whereas the latter four reconstruct the final result given a single image as input.
@@
4.1 Image-to-3D Garments
Given an image prompt and a base mesh, Garment3DGen generates textured 3D garments matching the input image guidance. We present a plethora of image-to-3D results in Fig. 1, Fig. 3 and Fig. 6 and showcase that the generated 3D assets can be of various topologies, respect the pose, shape and texture of the input image, garments can be both real and fantastical, contain high-quality texture maps, contain fine-level details such as cloth wrinkles and folds and preserve the topology and structure of the input (e.g., head, bottom and armholes are respected such that the garments can be draped on a body). In addition we quantitatively evaluate Garment3DGen against several baseline methods and report our results in Table 1. Our approach outperforms all methods in terms of both embedding as well as perceptual similarity with the input image guidance while it is the only approach that generates textured geometries that can be used for downstream tasks. Deformation-based approaches (rows 1-2) preserve the topology and the holes of the garment but lack strong supervision signals to generate outputs that match the input image. Reconstruction-based approaches (rows 3-6) are good at preserving and reconstructing what is visible in the input image but generate unusable geometries (or splats) and their color estimates for the non-visible regions are usually blurry or single colored (as shown in Fig. 6).
4.2 Text-to-3D Garments
Our work can be extended to generating textured 3D garments given textual prompt inputs. Unlike TextDeformer that utilizes a text-prompt we opt for a text-to-image diffusion model as an initial step because image-based supervisions can provide a stronger guidance for our mesh deformation process. Such an approach allows the user to iterate with several text-prompts to generate the image of their desired real or fantastical garment in contrast to providing a text-prompt and waiting for the 3D asset creation process to finish to see if the result matches their intent. In Fig. 3 (row 2) and Fig. 6 (rows 3-6) we present a variety of generated garments using only the provided textual prompt as guidance.
@@ 4.2 Text-to-3D Garments
4.3 Ablation Studies
@@
We conduct an ablation study to assess the impact of the key components of our proposed approach. We start from a basic baseline approach and add components one at a time and showcase our findings in Fig. 4. We start with the off-the-shelf TextDeformer [18] where given a text prompt and a base-mesh we deform the input geometry to match the target text using their framework and losses. Text prompts are not ideal to capture the fine-level details of a garment as there can be many “medieval armors” and in addition a pretrained MetaCLIP [74] model is not capable of capturing the subtle differences between a “jacket” and a “puffer jacket”. To overcome this limitation we adapt TextDeformer to take image inputs as guidance (ImageDeformer) and observe that the deformed geometries are one step closer to what we are after but still they fail to capture the details of the image. By swapping out the original MetaCLIP model and introducing a model fine-tuned on fashion data we observe that subtle details are better preserved across garments. An image-based reconstruction approach faithfully represents the geometry but generates coarse and watertight meshes that are unusable for downstream tasks. We use these meshes as pseudo ground-truth for our approach. Our Garment3DGen results in garments that faithfully follow the image guidance while containing the wrinkles and fine details. However the quality of the output geometry is not always ideal for physics-based downstream tasks (e.g., cloth simulation) which is why we introduced additional 3D supervisions that preserve a better mesh quality, see Fig. 5. All prior works either do not follow the input image guidance or generate low-quality geometries that cannot be used for downstream tasks.
Figure 5: Mesh Quality (Top) and Geometry Comparisons (Bottom): We showcase the wireframes of all approaches. Our method stands out as the only one that produces geometries that adhere to the input image while maintaining good mesh quality and incorporating necessary holes for physics-based simulation tasks. At the bottom we showcase the output geometry of various techniques to highlight that our approach captures fine geometric details without geometric artifacts (Wonder3D).
4.4 Applications & Discussion
Figure 6: Qualitative Comparisons: We demonstrate several mesh generation methods starting from an input image shown on the left. We show front and back views of each reconstruction. The 3D Gaussian Splatting [84] method generates distorted frontal colors and dark or blurry back colors while its geometry is not suitable for downstream tasks such as simulation. The second reconstruction approach [40, 83] generates watertight meshes with very coarse geometric details and blurred out colors. Our proposed approach outputs 3D geometries that are geometrically correct with fine-level texture details that prior works fail to generate.
Physics-based Cloth Simulation Automatically generated garments by Garment3DGen are simulation-ready and can be used directly in downstream applications, see Fig. 1, Fig. 7 and the supplemental material. We consider the generated 3D shape as rest shape for our simulations and incorporate a zero rest-angle dihedral energy to model the out-of-plane banding of the fabric. Fantastical garments such as armors are not well modeled using a zero rest-angle as they would lose their distinct shape during simulation and wrinkle unnaturally. To obtain visually pleasing results, we take the rest angle to be the 3D generated mesh one which allows it to maintain its shape throughout the simulation.
Hand-Cloth Interaction in VR Our garments are suitable for real-time simulations with hand-interaction using modern VR headsets. The rightmost part of Fig. 1 as well as the last row of Fig. 7 show a user interacting with the garments through real-time simulation with integrated hand-tracked interactions.
Garment Resizing Generated garments can further be manipulated to achieve distinct looks through garment resizing. Fig.1 shows both the generated dress on the left as well as a larger version of it on the right, draped on the same body.
Sketch to Garment Fig. 7 shows an application to generate 3D simulation-ready garments starting from a rough sketch shown in the top left. Given a sketch, we generate a realistic image using ControlNet [81] which then serves as the input to our method to generate the corresponding 3D asset. We show simulation results of this automatically generated dress.
[…]
Limitations: Garment3DGen handles a variety of garment types both real and fantastical. Due to the requirement of a template mesh, there is a limitation to what garments can be generated whilst still providing good distortion-free meshes. This can be mitigated by providing a more diverse template library. Our estimated textures, while faithful to the image, sometimes do not fully preserve fine-level details. We plan to address this by tuning the texture enhancement module to be conditioned on the reference image across all views while maintaining its multi-view color consistency properties.
Figure 7: Applications: Garment3DGen can generate textured 3D garments from images, from text prompts, from simple sketches, that can be fitted to human bodies and drive them with physics-based cloth simulation or even enable interaction between hands and garments in a VR environment.
5 Conclusion
We proposed Garment3DGen a new approach to generate high-quality, physically plausible garment assets that can be directly used for downstream applications. We introduced a deformation-based approach that takes a base mesh as input along with an image guidance and outputs a textured geometry that faithfully matches the input image while preserving the structure and topology of the input mesh. Our key contributions stem from utilizing novel diffusion-based generative models to synthesize 3D pseudo ground-truth that can be used as a soft supervision signal along with additional regularizations, a texture enhancement module that generates high-fidelity texture maps and a body-cloth optimization framework that fits the generated 3D garments to parameteric bodies. Our approach clearly outperforms prior work across all metrics while producing physically plausible and high-quality garment assets. Finally, we showcased several applications where the output geometries of Garment3DGen are used for physics-based cloth simulation, hand-garment interaction in a VR environment as well as going directly from a simple sketch image to a drivable 3D garment.

### v2

@@ 3.3 Texture Estimation
Given the untextured deformed 3D geometry $M_{\text{def}}$, we generate high-fidelity textures that match the input image. We leverage a 2D text-to-image generation model to create high-quality, high-resolution textures with vivid details given a text prompt. When starting from images where text is absent, a text prompt is obtained using an image caption model [40]. Our pipeline consists of the following steps: generating images of the given fashion asset from multiple views and backprojecting these images onto the mesh surface to create UV texture $T\in\mathbb{R}^{H\times W\times C}$. There are two challenges in adapting a 2D generation model to 3D objects: establishing geometry-texture correspondence and ensuring multi-view consistency.
@@
Multi-view Consistency: Directly conducting view-by-view image synthesis cannot guarantee that the generated views are consistent with each other. To solve this problem, we first synthesize the front and back views simultaneously to implicitly enforce global consistency. Then, we leverage the fact that the geometry of fashion assets is mostly flat, and conduct depth-aware inpainting for the remaining views to ensure the newly generated textures are locally consistent. To handle occluded areas, we design an automatic view selection algorithm to inpaint the textures in a coarse-to-fine manner: from the remaining unpainted area, we select the view with the most unfilled pixels to generate the textures iteratively from large regions to small pieces.
@@
Our texture estimation module has the following improvements compared with the previous works:
  • Global consistency: by generating the front and back views together, we enforce better global consistency.
  • Geometry-aware texture inpainting: Past works tweaked the depth-to-image generation model with masked generation methods (TEXTure [65] uses Blended Diffusion, and Text2Tex [10] uses RePaint). We directly use multi-ControlNet, enabling a better balance between the depth and mask controls, which produces better local content consistency.
  • Backprojection: we implement a faster backprojection method using PyTorch $grid\_sample$, which has the lowest latency compared to all known methods.
  • Speed: Our texture generation is significantly faster than: TEXTure: 2.5 mins, Text2Tex:12mins, Ours: 4.56s
We refer the reader to the supplementary for an in-depth discussion of trade-offs.
Figure 3: Qualitative Comparisons: We demonstrate several mesh generation methods given an image (top) or text (bottom) and show front and back views of each reconstruction. The 3D Gaussian Splatting [98] method generates distorted frontal colors and dark or blurry back colors while its geometry is not suitable for downstream tasks such as simulation. The second reconstruction approach [51, 97] generates watertight meshes with coarse geometric details and blurry colors. WordRobe[73] which is purely text-based, generates simulation-ready garments but they deviate far from the text prompt (e.g., the puffer jacket and t-shirt are not faithful to the prompt). Our proposed approach outputs 3D geometries that are geometrically correct with fine-level texture details that prior works fail to generate.
@@ 4 Experiments
Data: We use the publicly available DiffAvatar [43] dataset which comprises of artist-created garment templates covering several cloth categories (e.g. T-shirt, shirt, tank-top, dress, etc). To demonstrate garment generation using image guidance, we collect a variety of real images with garments in different poses, different textures, including garment types that are not covered by our dataset. In addition to real images, we include AI generated images using a text prompt of both realistic and fantastical garments.
@@
Metrics: Quantitatively evaluating our results is a challenging task in the absence of 2D or 3D ground-truth for what are we trying to accomplish. However we can evaluate how consistent the geometries are to the input image and hence we render untextured outputs for all methods from 36 views and compute their perceptual scores using the LPIPS metric [96] as well as their image-based CLIP similarity score using the cosine distance between their embeddings.
@@
Baselines: We evaluate our approach against: i) TextDeformer [24] which deforms input meshes based on text, ii) ImageDeformer - a variation of TextDeformer we developed where the input text is replaced with an image, iii) Wonder3D [51] and iv) Zero123++ [71] both of which generate 3D geometries given a single image using 2D diffusion models, v) ZeroShape [32] which performs zero-shot reconstruction and vi) a triplane-based 3DGS approach [98]. The first two works also take as input a base mesh whereas the latter four reconstruct the final result given a single image as input. Finally, we perform qualitative comparisons with WordRobe [73] which is a recent top-performing method among some concurrent text-based garment generation methods [31, 49, 39] and a pattern-based approach termed SewFormer [46] (see supplementary).
Figure 4: Mesh Quality (Top) and Geometry Comparisons (Bottom): We showcase the wireframes of all approaches. Our method stands out as the only one that produces geometries that adhere to the input image while maintaining good mesh quality and incorporating necessary holes for physics-based simulation tasks. At the bottom we showcase the output geometry of various techniques to highlight that our approach captures fine geometric details without geometric artifacts (Wonder3D). Note that several methods produce smooth meshes where details are incorrectly captured in the texture map instead which results in lower quality visual results.
4.1 Image-to-3D Garments
Given an image prompt and a base mesh, Garment3DGen generates textured 3D garments matching the input image guidance. We present a plethora of image-to-3D results in Fig. 1, Fig. 3 and Fig. 5 and showcase that the generated 3D assets can be of various topologies, respect the pose, shape and texture of the input image. Garments can be both real or fantastical. They contain high-quality texture maps which exhibits fine-level details such as cloth wrinkles and folds and preserves the topology and structure of the input (e.g., head, bottom and armholes are respected such that the garments can be draped on a body). Additionally, we quantitatively evaluate Garment3DGen against several baseline methods and report our results in Table 1. Our approach outperforms all methods in terms of both embedding as well as perceptual similarity with the input image guidance while it is the only approach that generates textured geometries that can be used for downstream tasks. Deformation-based approaches (rows 1-2) preserve the topology and the holes of the garment but lack strong supervision signals to generate outputs that match the input image. Reconstruction-based approaches (rows 3-6) are good at preserving and reconstructing what is visible in the input image but generate unusable geometries (or splats) and their color estimates for the non-visible regions are usually blurry or single colored (as shown in Fig. 3).
4.2 Text-to-3D Garments
We easily extend our work to enable textured garment generation using textual prompt inputs. Unlike TextDeformer [24] or WordRobe [73] which utilize text directly, we opt for a text-to-image diffusion model as an initial step. We do this because image-based supervisions provides a stronger guidance for our mesh deformation process which produces better results. This allows the user to easily iterate with several text-prompts to generate the image of their desired garment . This is in stark contrast to the slow iteration time of providing a text-prompt and waiting for the 3D asset creation process to finish to see if the result matches their intent. In Fig. 3 (bottom) we present generated garments using only the provided textual prompt as guidance.
@@ 4.2 Text-to-3D Garments
Figure 5: Applications: Garment3DGen generates textured 3D garments from images, text prompts, simple sketches, that can be fitted to human bodies and drive them with physics-based cloth simulation or even enable interaction between hands and garments in VR.
4.3 Ablation Studies
We conducted a variety of ablation studies to assess the impact of the key components of our proposed approach the figures of which are provided in the supplementary material. We start with the off-the-shelf TextDeformer which takes a text prompt and a base-mesh and deforms this to match the target text. Text prompts are not ideal to capture the fine-level details of a garment as there can be many “medieval armors”. In addition, a pretrained CLIP model is not capable of capturing the subtle differences between a “jacket” and a “puffer jacket”. To overcome this limitation we adapt TextDeformer to take image inputs as guidance (ImageDeformer) and observe that the deformed geometries are improved. Nonetheless, they still fail to capture the details of the image. By swapping the original CLIP model and introducing a model that is fine-tuned on fashion data we observe that details are better preserved across garments. Noting that image-based reconstruction methods can accurately capture geometry but produce coarse and watertight meshes that are unsuitable for subsequent tasks, we utilize these meshes as pseudo ground-truth for our proposed approach. Our Garment3DGen results in garments that faithfully follow the image guidance while containing the wrinkles and fine details. However the quality of the output geometry is not always ideal for physics-based downstream tasks because they produce poorly conditioned triangles which will result in instabilities when simulated in addition to poorly tessellated geometry which will result in unnatural fabric behavior. Because of this, we introduced additional 3D supervisions that preserve a better mesh quality, see Fig. 4. All prior works either do not follow the input image guidance or generate low-quality geometries that cannot be directly used for downstream tasks. Finally we conducted ablation studies to assess the impact of the pre-trained CLIP model on garment data, the impact of regularizations as well as the texture enhancement module and refer the interested reader in the supplementary material for additional information.
4.4 Applications
Physics-based Cloth Simulation We consider the generated 3D shape as rest shape for our simulations and incorporate a zero rest-angle dihedral energy to model the out-of-plane banding of the fabric. Fantastical garments such as armors are not well modeled using a zero rest-angle as they would lose their distinct shape during simulation and wrinkle unnaturally. To obtain visually pleasing results, we take the rest angle to be the 3D generated mesh one which allows it to maintain its shape throughout the simulation. Generated garments can further be manipulated to achieve distinct looks through garment resizing.
Hand-Cloth Interaction in VR Our garments are suitable for real-time simulations with hand-interaction using modern VR headsets. The rightmost part of Fig. 1 and Fig. 5 show a user interacting with the garments through real-time simulation with integrated hand-tracked interactions.
Sketch to Garment: Given a rough dress sketch, we generate a realistic image using ControlNet [95] which then serves as the input to our method to generate the corresponding 3D asset. In Fig. 5 we show simulation results of this automatically generated dress.
4.5 Discussion
Simulation-ready Garments: Our method stands out as the favorable choice for producing garments that are faithful to the prompt with high-quality meshes that can be seamlessly integrated into simulation frameworks as demonstrated by a qualitative and quantitative analysis. Although it might be possible to produce simulation-ready garments using alternative methods, this would require remeshing and potentially several additional manual post-processing steps, such as creating arm holes. Moreover, our use of template meshes ensures a superior UV layout, which is challenging to attain automatically.
[…]
Limitations: Garment3DGen handles a variety of garment types both realistic and fantastical. Due to the requirement of a template mesh, there is a limitation on what garments can be generated whilst still providing distortion-free meshes. This can be mitigated by providing a more diverse template library. Our estimated textures, while faithful to the image, sometimes do not fully preserve fine-level details. We plan to address this by tuning the texture enhancement module to be conditioned on the reference image across all views while maintaining its multi-view color consistency properties. Finally, it is worth noting that the closer the input template mesh is to the target garment, the easier the deformation task becomes. We currently select the closest template manually but performing automatic retrieval based on the image could be tackled in future work.
5 Conclusion
We proposed a new approach to generating high-quality garment assets that can be directly used for downstream applications which require good mesh quality. We introduced a deformation-based approach that takes a base mesh as input along with an image guidance and outputs textured geometry that faithfully matches the input image while preserving the structure and topology of the input mesh. Our key contributions stem from utilizing novel diffusion-based generative models to synthesize 3D pseudo ground-truth that can be used as a soft supervision signal along with additional regularizations, a texture enhancement module that generates high-fidelity texture maps and a body-cloth optimization framework that fits the generated 3D garments to parametric bodies. Our approach clearly outperforms prior work. Finally, we showcased physics-based cloth simulation, hand-garment interaction in a VR environment.
Acknowledgments
We thank Michelle Guo, Will Gao and Astitva Srivastava for their help and support on running baseline comparisons.
Supplementary Material
We refer the interested reader to the supplemental video where we provide a wide variety of results ranging from image/text to 3D textured garments as well as applications of our method in downstream tasks such as physics-based cloth simulation, hand-garment interaction in VR using a headset and sketch to 3D garment reconstruction. Below we provide some additional details regarding the implementation of our key components as well as some additional ablation studies to showcase the impact of our design decisions.
Garment3DGen General Details
We believe that our approach provides three key insights that will be valuable to the community:
  1. 1.
    Mesh-based deformations provide the right properties to generate (or stylize) new garments that we can utilize for downstream tasks other than rendering.
  2. 2.
    A text-prompt or a single image alone cannot provide enough guidance to generate the desired garment exactly the way a user might want it. This is evident from the results of WordRobe [73] which despite its mesh-quality the generated garments do not follow the provided text prompt.
  3. 3.
    3D supervisions, if done right, can provide strong enough supervision signal in order to generate the desired garments with the proper topology and structure.
Our approach builds upon these insights and introduces a novel yet simple solution to generate high-quality, physically plausible garments. As input to the method, we require only a single garment image (or alternatively, a text prompt that can generate this image using a text to image model) and a base garment template mesh. The input image needs to contain a single piece of clothing, captured from a semi-frontal viewpoint with its pose being as occlusion free as possible. A person can be wearing this garment or there might be more than one piece of clothing in the image in which case we perform semantic segmentation (using SAM) to obtain the garment. The template mesh is not required to be similar to the image guidance. For example, we demonstrate results where our method can go from a shirt to a puffer jacket, from a tank-top to a dress or even a T-shirt to a fantastical sea armor. Note that the closer the base mesh is to the target geometry, the easier the task is. For example, starting from a dress mesh to go to a shirt is a difficult task while starting from something closer to the target simplifies this problem.
Automatic View Selection: The goal of this algorithm is to automatically select the least-painted view and paint it. In this way, we can solve the 3D texture generation problem in a coarse-to-fine manner, and ensure the overall consistency. Alg. 1 provides a detailed description of the automatic view selection algorithm: given the input UV texture $T$ with painted front and back views, there could be $N$ candidate views. We maintain a binary mask $T^{B}$ that marks the painted pixel as 1, and unpainted pixel as 0. We can select the view with the most unfilled pixels as the next view to generate the appearance, and update the binary mask $T^{B}$. This process is repeated iteratively until most of the pixels are painted, or reaching a certain iteration number.
Input: an input mesh $M_{\text{def}}$ with UV texture $T$ with front and back views painted, a binary mask $T^{B}$ marking the painted pixels of $T$, and $N$ uniformly distributed candidate views $\{C_{i}\}_{i=1}^{N}$;
for number of iterations do
       Calculate the binary mask $T^{B}_{i}$ for each view $i$ from $T^{B}$: $\{T_{i}^{B}\}_{i=1}^{N}$;
       Select the least painted view $C_{j}$: $j\leftarrow\arg\min_{i=1}^{N}\sum T_{i}^{B}$ ;
       Generate the appearance image $I_{i}$ and update $T^{B}$;
end for
Algorithm 1 Automatic View Selection
Table 2: Comparisons of different texture estimation methods. The runtime is measured on a single NVIDIA H100 GPU.
Method Pros Cons Runtime
Mesh2Tex [6] Infinite resolution & Global consistency One model per class & No fine details 8mins
TEXTure [65] Shape-aware & Local consistency Bad Global consistency & texture artifacts & Janus 2mins
Text2Tex  [10] Shape-aware & Local consistency Bad Global consistency & color/pattern shifts & Janus 5mins
[…]
Mesh Deformer Details
  • Alignment: Using the nvdiffmodelling library the base mesh is aligned to the target mesh using the unit-size function that moves/rescales the input to match bounding boxes.
  • Deformation: We use the same formulation with the Neural Jacobian Fields(NJF) as described in Sec.3.1 of their paper and Sec.3 of TextDeformer(TD). Once the deformation map is obtained using Eq.(1) we obtain the updated vertices of the input mesh. While NJF could deform a garment to have a different pose it’s not possible to change its style (t-shirt-¿sea-creature-armor) because of the supervision signals. Our goal was not to do mesh-registration but instead stylize input base garment templates via deformation. Hence NJF was chosen for its versatility across heterogeneous mesh collections, because it’s triangulation-agnostic and it provided a flexible and easy-to-use framework to accomplish our goal in a fast plug-n-play manner.
  • Losses: Our goal was to deform the base mesh enough to match the pseudo-ground-truth mesh extracted from the image but not fully since if that was the case we’d end up with a watertight mesh unable to be fit to parametric bodies and simulated. Hence we opted for point-to-point meshes (instead of point-to-mesh) for the 3D supervision as well as embedding and image-based losses in the 2D space.
Texture Details
This approach prioritized filling in the large areas first before moving on to smaller and more occluded regions.
Texture Comparisons: In Table 2 we provide a comparison between the pros and cons of recent texture estimation approaches. Our approach is significantly faster compared to past works due to key optimizations described in the main paper while maintaining good local and global consistency. Similar to past works one can notice the Janus problem appearing in texture maps, which can be handled by training a multiview generation diffusion model with more explicit camera pose injection in the future works.
Figure 6: Fitting a parametric body to a generated garment: We start with the generated textured 3D garment (in this case a medieval armor) and a parametric body in its canonical pose (left). After the body-garment optimization process the body pose and shape parameters are optimized such that the generated garment can fit in the body accurately without penetrations.
Quantitative Comparisons - Details: In terms of garment base meshes we utilized the publicly available dataset provided by DiffAvatar which comprises 6 template geometries. For each one of the 6 garments we provide 4 different image inputs (2 real and 2 AI-generated) and to quantitatively evaluate the different approaches we render the untextured outputs of all methods from 36 views. No prompts are used during the quantitative evaluation procedure.
Garment Fitting to Parametric Bodies
While the aforementioned supervisions and regularizations aim to ensure that the quality of the generated garments will be satisfactory for simulation. The produced garment will still need to be scaled, positioned and oriented to fit the parametric body [52] on which it is to be draped and simulated. To accomplish this task, we run an optimization procedure during which the generated garment remains fixed in the generated pose and the pose and shape of the parametric body are transformed such that the garment can accurately fit the body. This optimization process shown in in Fig. 6, starts with a rigid transformation and scaling of the body and continues with an optimization of the body pose and shape using the Chamfer distance loss.
Rigid Transformation and Scaling: The optimization is initialized by applying a rigid transformation (rotation, translation and scaling) of the parametric body model. This step roughly aligns the body with the garment.
Figure 7: Impact of the pre-trained CLIP on garment data: We disable all other supervisions and explore the impact of a pre-trained CLIP model on fashion data versus using the regular model to enforce embedding supervisions. We observe that regular CLIP embeddings result in distorted and unusable geometries regardless of whether the input is a text prompt or an image.
Pose and Shape Optimization: Subsequently, the body pose and shape parameters are optimized to minimize the Chamfer distance between the body mesh and the garment mesh.
Collision Handling: After body model optimization, an additional step is performed to resolve potential body-cloth collisions. This is achieved by minimizing an interpenetration loss that penalizes any intersections between the body mesh and the garment mesh.By combining these steps, the garment mesh can be accurately fit to the parametric body model, enabling realistic draping and physics-based cloth simulations for downstream applications.
Figure 8: Impact of regularizations on the final armor geometry: Enforcing no regularizations (Laplacian smoothing, penalization of small triangles etc.) on the output mesh results in a crisp output armor mesh with arm/body holes but its quality is not at the level required to perform physics-based simulation. On the other hand, enforcing strong regularizations results in overly smoothed meshes with closed holes. Our output strikes a good balance between capturing those fine-level details that make an armor geometry look like one yet making it suitable for downstream tasks.
Additional Ablation Studies
[…]
Figure 9: Impact of Texture Module: given the left image as a condition, the texture enhancement module enriches the details and enhances the overall image quality by effectively utilizing the powerful 2D priors.
Figure 10: 3D Garment Generation: Given an image ($1^{st}$ row) or a text prompt ($2^{nd}$ row) as guidance and a base geometry mesh (bottom left inset) that can be far from the target we generate high-quality textured 3D geometries of both real as well as fantastical garments.
@@ Acknowledgments
Figure 12: Qualitative Comparisons: SewFormer [46] produces 2D garment patterns from a single RGB image input. We further process these panels by virtually sewing them together at the seams to create assembled 3D garments, which enables us to perform a qualitative comparison. We showcase two garment examples and demonstrate that Garment3DGen produces results that are more faithful to the image input with a less complex pipeline.
@@
Adding Components one at a time: As described in the main paper we conducted an ablation study depicted in Fig. 11. We start with the off-the-shelf TextDeformer which takes a text prompt and a base-mesh and deforms this to match the target text. Text prompts are not ideal to capture the fine-level details of a garment as there can be many “medieval armors”. In addition, a pretrained CLIP model is not capable of capturing the subtle differences between a “jacket” and a “puffer jacket”. To overcome this limitation we adapt TextDeformer to take image inputs as guidance (ImageDeformer) and observe that the deformed geometries are improved. Nonetheless, they still fail to capture the details of the image. By swapping the original CLIP model and introducing a model that is fine-tuned on fashion data we observe that details are better preserved across garments. Noting that image-based reconstruction methods can accurately capture geometry but produce coarse and watertight meshes that are unsuitable for subsequent tasks, we utilize these meshes as pseudo ground-truth for our proposed approach. Our Garment3DGen results in garments that faithfully follow the image guidance while containing the wrinkles and fine details. However the quality of the output geometry is not always ideal for physics-based downstream tasks because they produce poorly conditioned triangles which will result in instabilities when simulated in addition to poorly tessellated geometry which will result in unnatural fabric behavior. Because of this, we introduced additional 3D supervisions that preserve a better mesh quality.
Texture Module: The impact of the texture enhancement module is shown in Fig. 9. The textures directly synthesized by 3D generation models tend to be low-resolution, smooth and over-simplified, which is due to the scarcity of high quality 3D training data. Thus, the texture enhancement module aims to effectively utilize the 2D priors learned from the large high-quality image dataset. After our image-conditioned image enhancement, we bring back vivid details to the texture, improving the perceptual quality.
Additional Results
Comparisons with SewFormer: In Fig. 12 we provide a qualitative comparisons against SewFormer [46] which predicts garment patterns from a single RGB image.To facilitate a meaningful comparison, we use a physics-based cloth simulator to sew the generated panels together, creating an assembled 3D garment. In the top row T-shirt example, SewFormer incorrectly generates pattern pieces for a dress instead of the requested T-shirt. After assembly, it’s clear that the predicted garment resembles a dress rather than the intended T-shirt. The second example shows that while SewFormer works reasonably well for certain garments, it fails to fully match the guidance, missing sleeves in this case. Both examples demonstrate that SewFormer produces results that don’t accurately match the image guidance, and its method for obtaining 3D assembled garments is more complex, requiring specialized software to sew individual pattern pieces together. Additionally, SewFormer is limited to producing V-neck designs, whereas our method correctly follows the visual guidance.
Additional Qualitative Comparisons: In Fig. 10 we provide multi-view renders of our 3D textured garments that we generate from text prompts and an image guidance. From these results we gain the following insights: a) Garment3DGen works just as well with fantastical garments (armors or dresses) that are outside the regular garment distribution, b) our texture estimation module results into high-quality textures that closely match the input text prompt and c) our output geometry does not have to be similar to the input base mesh. Finally in Fig. 13 we showcase the plethora of applications that Garment3DGen has ranging from text/image/sketch to simulation-ready 3D garments to hand-garment interaction in a VR environment using the on-device hand-tracking.
Figure 13: Garment3DGen Applications: We showcase various applications of our proposed approach.